What quantity of a 55% acid solution must be mixed with a 25% solution to produce 400 mL of a 40% solution?
Found 2 solutions by Alan3354, Maths68:
Answer by Alan3354(69443)

What quantity of a 55% acid solution must be mixed with a 25% solution to produce 400 mL of a 40% solution?
--------------
40% is the average of 55% and 25%, so it's equal amounts.

Answer by Maths68(1474)

Solution A
Amount = x
Concentration = 55%=0.55
Solution B
Amount = 400-x
Concentration = 25%=0.25
Resultant Solution
Amount = 400
Concentration = 40% =0.4


(Amount of A)(concentration of A)+ (Amount of B)(concentration of B)= (Amount of resultant)(concentration of resultant)
(x)(0.55)+(400-x)(0.25)=(400)(0.4)
0.55x+100-0.25x=160
0.55x-0.25x=160-100
0.3x=60
0.3x/0.3=60/0.3
x=200

200ml of a 55% acid solution must be mixed with a 200ml of 25% solution to produce 400 mL of a 40% solution